As performed in the ELOG 202, the PZTs of the OMC 002 were tested.
DC response was measured by sweeping each PZT with 0-150V triangular voltage at 11Hz. Acquire 0.2sec of the tie series using an oscilloscope to get the PDH error, cavity transmission, and the sweep signal.
The voltage where the tranmission peaks were observed were fitted were recorded. One fringe corresponds to the displacement of 532nm. So the displacement and the applied volatagewere fitted witha linear function.
This gave the PZT response for PZT1 and PZT2 to be 14.9nm/V and 14.4nm/V.
AC response was measured with SR785. The PZT was shaken with 1~50mVpp signal with the DC offset of 5V while the OMC was locked with the feedback to the laser fast PZT. The transfer function from the applied PZT voltage to the servo output were measured. The closed loop TF was also measured to remove the effect of the servo control. The DC levels of the responses were calibrated using the values above.